# StoxLSTM: A Stochastic Extended Long Short-Term Memory Network for Time Series Forecasting

### v1

@@ 3.2.1 Design of the Recurrent Unit in StoxLSTM
By incorporating stochastic latent variables $z_{1:t}$ into the original xLSTM recurrent unit, as shown in Figs. 1 and 2, StoxLSTM follows the SSM that respects temporal causality, Markovianity, and non-autoregressive dynamics. Specifically, at each time step $t$, the latent state $z_{t}$ is modeled as a stochastic variable conditioned on the previous latent state $z_{t-1}$ and the current deterministic hidden state $h_{t}$.
@@
Firstly, StosLSTM (the stochasticized sLSTM block in xLSTM, as shown in Fig. 1(a)) forward pass is:
@@
$\begin{array}[]{ll}{c_{t}}={f_{t}}^{\prime}{c_{t-1}}+{i_{t}}^{\prime}{y_{t}}&\text{cell state}\\ {n_{t}}={f_{t}}^{\prime}{n_{t-1}}+{i_{t}}^{\prime}&\text{normalizer state}\\ {h_{t}}={o_{t}}\widetilde{{h_{t}}}&\text{hidden state}\\ z_{t}\sim\mathcal{N}\big{(}\mu_{\theta}(h_{t},z_{t-1}),\sigma_{\theta}(h_{t},z_{t-1})\big{)}&\text{latent state}\\ {x_{t}}=\varphi\left({w_{z}^{T}{z_{t}}+{r_{z}}{h_{t}}+{b_{z}}}\right)&\text{output gate}\end{array}$
Here, $f_{t}$$^{\prime}$ and $i_{t}$$^{\prime}$ denote the stabilized forget gate and stabilized input gate, respectively. $y_{t}$} represents the cell input, and $o_{t}$ denotes the output gate.
@@
Secondly, StomLSTM (the stochasticized mLSTM block in xLSTM, as shown in Fig. 1(b)) forward pass is:
@@
$\begin{array}[]{ll}{c_{t}}={f_{t}}{c_{t-1}}+{i_{t}}{v_{t}}k_{t}^{T}&\text{cell state}\\ {n_{t}}={f_{t}}{n_{t-1}}+{i_{t}}{k_{t}}&\text{normalizer state}\\ {h_{t}}={o_{t}}\odot{\rm{}}{\widetilde{h}_{t}},\quad{\widetilde{h}_{t}}={C_{t}}{q_{t}}/max\left\{{\left|{n_{t}^{T}{q_{t}}}\right|,1}\right\}&\text{hidden state}\\ z_{t}\sim\mathcal{N}\big{(}\mu_{\theta}(h_{t},z_{t-1}),\sigma_{\theta}(h_{t},z_{t-1})\big{)}&\text{latent state}\\ {x_{t}}=\varphi\left({w_{z}^{T}{z_{t}}+{r_{z}}{h_{t}}+{b_{z}}}\right)&\text{output gate}\end{array}$
@@
Here, $k_{t}$ denotes the key input, $q_{t}$ denotes the query input, and $\odot$ represents element-wise multiplication.
@@
In both StosLSTM and StomLSTM, we assume that the latent state $z_{t}$ follows a Gaussian distribution parameterized by a mean $\mu_{\theta}(h_{t},z_{t-1})$ and variance $\sigma_{\theta}(h_{t},z_{t-1})$, both predicted by neural networks. This probabilistic formulation allows the model to capture the inherent uncertainty and stochasticity present in real-world time series data. By sampling $z_{t}$ from this distribution through the reparameterization block, StoxLSTM is able to represent diverse possible future trajectories, effectively capturing the variability and noise inherent in temporal dynamics. The update of the latent state $z_{t}$ thus follows the conditional distribution $p(z_{t}\mid z_{t-1},x_{1:t-1})$. Similarly, the output $x_{t}$ is generated by applying a linear transformation to $z_{t}$ and $h_{t}$, followed by an activation function $\varphi$. The generation of $x_{t}$ follows the conditional distribution $p(x_{t}\mid z_{t},x_{1:t-1})$.
@@ 3.2.2 Generative Model
The generative model evolves according to the factorization in Eq. (1). During preprocessing, zero-padding is applied at both ends of the look-back sequence: padding at the front generates the initial input $x_{0}$ with a length equal to the patch stride $S$, and padding at the end aligns the predicted sequence length $T$. The padded sequence is then segmented into patches $p_{0:N}$, where $N=\lceil\frac{L+T+S-P}{S}\rceil$, with patch size $P$. These patches are embedded as:
${x_{p0:pN}}={\rm{Embedding}}({\rm{patching}}([\underbrace{0,\ldots,0,}_{{S}}{x_{1:L}},\underbrace{0,\ldots,0}_{T}]))$ (3)
The embedded patches $x_{p0:pN}$ fed into the StoxLSTM recurrent unit to generate latent states $z_{1:N+1}$ and time series patches $x_{p1:pN+1}$:
$z_{1:N+1},x_{p1:pN+1}=\mathrm{RecurrentUnit}(x_{p0:pN})$ (4)
Therefore, during the reconstruction phase, the latent states $z_{t}$ and observations $x_{t}$ are generated from the probability distributions ${p_{\theta}}(z_{t}|z_{t-1},x_{1:t-1})$ and ${p_{\theta}}(x_{t}|z_{t},x_{1:t-1})$, respectively. While during the forecasting phase, $z_{t}$ and $x_{t}$ are generated from the distributions ${p_{\theta}}(z_{t}|z_{t-1},x_{1:L})$ and ${p_{\theta}}(x_{t}|z_{t},x_{1:L})$, respectively.
The time series patches $x_{p1:pN+1}$ are then flattened and passed through a linear layer to generate the final prediction $x_{L+1:L+T}$:
$x_{L+1:L+T}=\mathrm{FCL}(\mathrm{Flatten}(x_{p1:pN+1}))$ (5)
@@ 3.2.3 Inference Model
The inference model follows the factorization in Eq. (2) via a bidirectional recurrent unit [40, 19]. Zero-padding is applied at the front of the sequence $x_{1:L+T}$ with a length equal to the patch stride, followed by patching and embedding:
${x_{p0:pN}}={\rm{Embedding}}({\rm{patching}}([\underbrace{0,\ldots,0,}_{\rm{S}}{x_{1:L+T}}]))$ (6)
The bidirectional recurrent unit processes these embedded patches, producing forward and backward hidden states $h_{1:N+1}$ and $g_{N+1:1}$, which are then used to generate the latent states $z_{1:N+1}$:
${z_{1:N+1}},\_\ =\mathrm{RecurrentUnit_{bidirection}}(x_{p0:pN})$ (7)
The bidirectional recurrent unit enables the latent states $z_{t}$ to be generated from the approximate posterior distribution ${q_{\varphi}}\left({{z_{t}}|{z_{t-1}},{x_{1:L+T}}}\right)$. The role of the inference model is to obtain this approximate posterior during training, while during testing, only the generative model is utilized.
@@ 3.2.4 Evidence Lower Bound of StoxLSTM
Here, $KL(\cdot)$ denotes the KL divergence. This objective encourages the model to reconstruct the observed sequence accurately while maintaining consistency between the inferred latent states and their temporal dynamics. Detailed derivations and implementation details are provided in Appendix A.
3.2.5 Complexity Analysis of StoxLSTM
For a time series $x_{1:L+T}^{C}$, the time complexity of xLSTM is $O(L+T)$, and its space complexity is $O(1)$ [14]. Considering that StoxLSTM adopts patching and channel independence, its time complexity is $O(C(L+T))$ and the space complexity is $O(1)$. Compared to Transformer-based models, StoxLSTM exhibits lower time and space complexity. However, it still has a relatively large number of parameters. During training, both the inference and generative models are required, but during inference, only the generative model is used, reducing computational time and memory overhead.
@@ 4.1.1 Datasets
We employ 9 publicly available datasets to evaluate the performance of our proposed StoxLSTM for long-term time series forecasting. These datasets include: Electricity [^1] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014, Weather [^2] https://www.bgc-jena.mpg.de/wetter/, Solar [^3] https://github.com/laiguokun/multivariate-time-series-data, Traffic [^4] http://pems.dot.ca.gov, ILI [^5] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html, and four ETT [^6] https://github.com/zhouhaoyi/ETDataset datasets (ETTm1, ETTm2, ETTh1, ETTh2). Additionally, we utilize 4 public datasets to assess the performance of short-term time series forecasting, which include: PEMS [^7] http://pems.dot.ca.gov (PEMS03, PEMS04, PEMS07, and PEMS08). These datasets have been widely used for benchmarking and are publicly available in [24], [27], [16] and [43]. Detailed dataset information is summarized in Table 1.
@@
Datasets Dim Horizon Dataset size Frequency Information Tasks
Weather 21 {96, 192, 336, 720} (36887, 5270, 10539) 10 min Weather Long-term forecasting
Electricity 321 {96, 192, 336, 720} (18412, 2632, 5260) Hourly Electricity Long-term forecasting
Solar-Energy 137 {96, 192, 336, 720} (36792, 5256, 10512) 10 min Electricity Long-term forecasting
ETTh1, ETTh2 7 {96, 192, 336, 720} (8640, 2880, 2880) Hourly Temperature Long-term forecasting
ETTm1, ETTm2 7 {96, 192, 336, 720} (34560, 11520, 11520) 15 min Temperature Long-term forecasting
Traffic 862 {96, 192, 336, 720} (12280, 1756, 3508) 10 min Transportation Long-term forecasting
ILI 170 {24, 36, 48, 60} (676, 97, 193) 7 days Illness Long-term forecasting
PEMS03 358 12 (15724, 5241, 5243) 5 min Transportation Short-term forecasting
PEMS04 307 12 (10195, 3398, 3399) 5 min Transportation Short-term forecasting
PEMS07 883 12 (16934, 5644, 5646) 5 min Transportation Short-term forecasting
PEMS08 170 12 (10713, 3571, 3572) 5 min Transportation Short-term forecasting
4.1.2 StoxLSTM Configuration
StoxLSTM is configured with a consistent set of hyperparameters across all experiments: Patch size $P=56$, Stride $S=24$, model dimension d_model=64d\_model=64, and latent dimension d_latent=16d\_latent=16. Other hyperparameters, including fully connected layer sizes, batch size, learning rate, and dropout rate, are adjusted based on dataset dimensionality and size.
In all experiments, we use mean squared error (MSE) as the reconstruction loss. The loss function of StoxLSTM can be expressed as:
${Loss}\_{{function}}=\text{MSE}\left(x_{1:L+T},\hat{x}_{1:L+T}\right)+\frac{\beta}{d}\sum_{i=1}^{d}\left(\log\frac{\sigma_{p_{i}}}{\sigma_{q_{i}}}-\frac{1}{2}+\frac{\sigma_{p_{i}}^{2}+(\mu_{q_{i}}-\mu_{p_{i}})^{2}}{2\sigma_{q_{i}}^{2}}\right)$
where $x_{1:L+T}$ denotes the ground truth sequence, and $\hat{x}_{1:L+T}$ represents the prediction from StoxLSTM. The second term corresponds to the KL divergence between two Gaussian distributions. Here, $p$ denotes either $p_{\theta}\left(z_{t}|z_{t-1},x_{1:L}\right)$ or $p_{\theta}\left(z_{t}|z_{t-1},x_{1:t-1}\right)$, while $q$ denotes $q_{\varphi}\left(z_{t}|z_{t-1},x_{1:L+T}\right)$. $\mu$ and $\sigma$ represent the mean and standard deviation of the respective distributions. $d$ is the latent dimension d_latentd\_latent, and $i$ indexes the $i$-th dimension of the latent variable. The hyperparameter $\beta$ controls the relative weighting between the reconstruction and KL divergence terms, and is typically set to $\beta=500$ in our experiments. The RAdam optimizer [44] is employed alongside a cosine annealing learning rate scheduler.
@@ 4.1.2 StoxLSTM Configuration
Transformer-based: Informer [13], Autoformer [24], Crossformer [25], FEDformer [26], PatchTST [32] and iTransformer [31].
@@
xLSTM-based: xLSTMTime [17], P-sLSTM [15] and xLSTM-Mixer [16].
@@
For long-term forecasting, the prediction horizons for the ILI dataset are set to $T\in\{24,36,48,60\}$, with StoxLSTM’s look-back window fixed at $L=96$. For all other datasets, the prediction horizons are $T\in\{96,192,336,720\}$, and the look-back window of StoxLSTM is set to $L=336$. For short-term forecasting, all models are set to $T=12$ and $L=96$. Baseline results and look-back length $L$ settings of long-term forecasting are primarily sourced from PatchTST [32], xLSTM-Mixer [16] and TimeMixer++ [29], supplemented by our own experiments where results are unavailable. Evaluation metrics for long-term forecasting include Mean Squared Error (MSE) and Mean Absolute Error (MAE), while short-term forecasting is assessed using MAE, Mean Absolute Percentage Error (MAPE), and Root Mean Squared Error (RMSE). Experiments are conducted on an NVIDIA RTX 3090 GPU (24GB) and 3 NVIDIA A800 GPUs (80GB).
4.1.4 Metrics
To ensure maximum comparability with prior work [32, 16, 29], we adopt the commonly used evaluation metrics for all models. Specifically, long-term forecasting is assessed using mean absolute error (MAE) and MSE. Short-term forecasting is evaluated with MAE, root mean squared error (RMSE), and mean absolute percentage error (MAPE). These metrics are averaged across all variables and calculated as follows:
${\rm{MAE}}\left({\textbf{y},\widehat{\textbf{y}}}\right)=\frac{1}{{{T}}}\sum\limits_{i=1}^{T}{\left|{{y_{i}}-{{\hat{y}}_{i}}}\right|}\qquad{\rm{MAPE}}\left({\textbf{y},\widehat{\textbf{y}}}\right)=\frac{{100}}{{{T}}}\sum\limits_{i=1}^{T}{\frac{{\left|{{y_{i}}-{{\hat{y}}_{i}}}\right|}}{{\left|{{y_{i}}}\right|+\varepsilon}}}$
${\rm{MSE}}\left({\textbf{y},\widehat{\textbf{y}}}\right)=\frac{1}{{{T}}}\sum\limits_{i=1}^{T}{{{\left({{y_{i}}-{{\widehat{y}}_{i}}}\right)}^{2}}}\qquad{\rm{RMSE}}\left({\textbf{y},\widehat{\textbf{y}}}\right)=\sqrt{\frac{1}{{{T}}}\sum\limits_{i=1}^{T}{{{\left({{y_{i}}-{{\widehat{y}}_{i}}}\right)}^{2}}}}$
wherein y denotes the ground truth, $\widehat{\textbf{y}}$ denotes the predictions, and $\varepsilon$ is a small constant for numerical stability. Additionally, the Continuous Ranked Probability Score (CRPS) evaluation metric is detailed in Appendix B.1.

### v2

@@ 3.2.1 Design of the Recurrent Unit in StoxLSTM
By incorporating stochastic latent variables $\bm{z}_{1:t}$ into the original xLSTM recurrent unit, StoxLSTM follows the SSMs that respect temporal causality, Markovianity, and non-autoregressive dynamics, as shown in Figs. 1 and 2. Specifically, at each time step $t$, the latent state $\bm{z}_{t}$ is modeled as a stochastic variable conditioned on the previous latent state $\bm{z}_{t-1}$ and the current deterministic hidden state $\bm{h}_{t}$.
@@
The forward pass of StosLSTM (the stochasticized sLSTM block in xLSTM, as shown in Fig. 1(a)) is as follows:
@@
$\begin{array}[]{ll}{\bm{c}_{t}}={\bm{f}_{t}}^{\prime}{\bm{c}_{t-1}}+{\bm{i}_{t}}^{\prime}{\bm{y}_{t}}&\text{cell state}\\ {\bm{n}_{t}}={\bm{f}_{t}}^{\prime}{\bm{n}_{t-1}}+{\bm{i}_{t}}^{\prime}&\text{normalizer state}\\ {\bm{h}_{t}}={\bm{o}_{t}}{\bm{\widetilde{h}}_{t}},\quad{\bm{\widetilde{h}}_{t}}={\bm{c}_{t}}/{\bm{n}_{t}}&\text{hidden state}\\ \bm{z}_{t}\sim\mathcal{N}\big(\mu_{\bm{\theta}}(\bm{h}_{t},\bm{z}_{t-1}),\sigma_{\bm{\theta}}(\bm{h}_{t},\bm{z}_{t-1})\big)&\text{latent state}\\ {x_{t}}=\varphi\left({\bm{w_{z}}^{T}{\bm{z}_{t}}+\bm{r_{z}}{\bm{h}_{t}}+\bm{b_{z}}}\right)&\text{output generation}\end{array}$
Here, $\bm{y}_{t}$ denotes the cell input. The variables $\bm{c}_{t}$, $\bm{n}_{t}$, and $\bm{h}_{t}$ represent the cell state, normalizer state, and hidden state, respectively. The latent state (latent variable) $\bm{z}_{t}$ follows a Gaussian distribution parameterized by mean $\mu_{\bm{\theta}}(\cdot)$ and standard deviation $\sigma_{\bm{\theta}}(\cdot)$, both implemented as fully connected neural networks. The terms $\bm{f}_{t}$$^{\prime}$ and $\bm{i}_{t}$$^{\prime}$ correspond to the stabilizer’s forget and input gates [3], while $\bm{o}_{t}$ denotes the output gate. Parameters $\bm{w_{z}}$, $\bm{r_{z}}$, and $\bm{b_{z}}$ are the weights and bias associated with $\bm{z}_{t}$ and $\bm{h}_{t}$. The function $\varphi$ represents the activation function. Finally, $x_{t}$ denotes the generated output (observation) at time step $t$.
@@
Similarly, the forward pass of StomLSTM (the stochasticized mLSTM block in xLSTM, as shown in Fig. 1(b)) is:
@@
$\begin{array}[]{ll}{\bm{c}_{t}}={\bm{f}_{t}}{\bm{c}_{t-1}}+{\bm{i}_{t}}{\bm{v}_{t}}\bm{k}_{t}^{T}&\text{cell state}\\ {\bm{n}_{t}}={\bm{f}_{t}}{\bm{n}_{t-1}}+{\bm{i}_{t}}{\bm{k}_{t}}&\text{normalizer state}\\ {\bm{h}_{t}}={\bm{o}_{t}}\odot\bm{\widetilde{h}_{t}},\quad\bm{\widetilde{h}_{t}}=\bm{C_{t}}{\bm{q}_{t}}/max\left\{{\left|{\bm{n}_{t}^{T}{\bm{q}_{t}}}\right|,1}\right\}&\text{hidden state}\\ \bm{z}_{t}\sim\mathcal{N}\big(\mu_{\bm{\theta}}(\bm{h}_{t},\bm{z}_{t-1}),\sigma_{\bm{\theta}}(\bm{h}_{t},\bm{z}_{t-1})\big)&\text{latent state}\\ {x_{t}}=\varphi\left({\bm{w_{z}}^{T}{z_{t}}+\bm{r_{z}}{\bm{h}_{t}}+\bm{b_{z}}}\right)&\text{output generation}\end{array}$
@@
Here, $\bm{k}_{t}$ and $\bm{q}_{t}$ denote the key and query inputs, respectively, and $\odot$ represents element-wise multiplication.
@@
In both StosLSTM and StomLSTM, the latent state $\bm{z}_{t}$ is modeled as a Gaussian random variable conditioned on $\bm{h}_{t}$ and $\bm{z}_{t-1}$. This probabilistic formulation enables the model to capture inherent uncertainty and stochasticity in real-world time series. By sampling $\bm{z}_{t}$ via the reparameterization block, StoxLSTM can represent diverse possible future trajectories, effectively modeling the variability and noise intrinsic to temporal dynamics. Consequently, the latent state update follows the conditional distribution $p(\bm{z}_{t}\mid\bm{z}_{t-1},x_{1:t-1})$, while the output $x_{t}$ is generated according to $p(x_{t}\mid\bm{z}_{t},x_{1:t-1})$.
@@ 3.2.2 Generative Model
The generative process adheres to the factorization presented in Eq. (1). During preprocessing, zero-padding is applied at both ends of the look-back sequence: padding at the front generates the initial input $x_{0}$ with a length equal to the patch stride $S$, and padding at the end aligns the predicted sequence length $T$. The padded sequence is then segmented into patches $\bm{p}_{0:N}$, where $N=\lceil\frac{L+T+S-P}{S}\rceil$ and $P$ denotes the patch size. These patches are embedded as follows:
${\bm{x}_{p0:pN}}={\rm{Embedding}}({\rm{patching}}([\underbrace{0,\ldots,0,}_{{S}}{x_{1:L}},\underbrace{0,\ldots,0}_{T}]))$ (3)
The embedded patches $\bm{x}_{p0:pN}$ are fed into the StoxLSTM recurrent unit to generate latent states $\bm{z}_{1:N+1}$ and time series patches $\bm{x}_{p1:pN+1}$:
$\bm{z}_{1:N+1},\bm{x}_{p1:pN+1}=\mathrm{RecurrentUnit}(\bm{x}_{p0:pN})$ (4)
Therefore, during the reconstruction phase, the latent state $\bm{z}_{t}$ and observation $x_{t}$ are generated from the probability distributions ${p_{\bm{\theta}}}(\bm{z}_{t}|\bm{z}_{t-1},x_{1:t-1})$ and ${p_{\bm{\theta}}}(x_{t}|\bm{z}_{t},x_{1:t-1})$, respectively. While during the forecasting phase, $\bm{z}_{t}$ and $x_{t}$ are generated from the distributions ${p_{\bm{\theta}}}(\bm{z}_{t}|\bm{z}_{t-1},x_{1:L})$ and ${p_{\bm{\theta}}}(x_{t}|\bm{z}_{t},x_{1:L})$, respectively.
Finally, The time series patches $\bm{x}_{p1:pN+1}$ are flattened and passed through a fully connected layer to generate the final prediction $x_{L+1:L+T}$:
$x_{L+1:L+T}=\mathrm{FCL}(\mathrm{Flatten}(\bm{x}_{p1:pN+1}))$ (5)
@@ 3.2.3 Inference Model
The inference model follows the factorization in Eq. (2) and employs a bidirectional recurrent unit [18, 13]. Zero-padding is applied at the beginning of the sequence $x_{1:L+T}$ with length equal to the patch stride $S$, followed by patching and embedding:
${\bm{x}_{p0:pN}}={\rm{Embedding}}({\rm{patching}}([\underbrace{0,\ldots,0,}_{\rm{S}}{x_{1:L+T}}]))$ (6)
The bidirectional recurrent unit processes these embedded patches, producing forward and backward hidden states $\bm{h}_{1:N+1}$ and $\bm{g}_{N+1:1}$, which are then used to infer the latent states $\bm{z}_{1:N+1}$:
${\bm{z}_{1:N+1}},\_\ =\mathrm{RecurrentUnit_{bidirection}}(\bm{x}_{p0:pN})$ (7)
The bidirectional recurrent unit enables the latent states $\bm{z}_{t}$ to be drawn from the approximate posterior distribution ${q_{\bm{\varphi}}}\left({{\bm{z}_{t}}|{\bm{z}_{t-1}},{x_{1:L+T}}}\right)$. The role of the inference model is to obtain this approximate posterior during training, while during testing, only the generative model is utilized.
@@ 3.2.4 Evidence Lower Bound of StoxLSTM
Here, $KL(\cdot)$ denotes the KL divergence. This objective encourages the model to accurately reconstruct the observed sequence while maintaining consistency between the inferred latent states and their temporal dynamics. Detailed derivations and implementation details are provided in A.
@@ 4.1.1 Datasets
We employ 9 publicly available datasets to evaluate the performance of our proposed StoxLSTM for long-term time series forecasting. These datasets include: Electricity [^1] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014, Weather [^2] https://www.bgc-jena.mpg.de/wetter/, Solar [^3] https://github.com/laiguokun/multivariate-time-series-data, Traffic [^4] http://pems.dot.ca.gov, ILI [^5] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html, and four ETT [^6] https://github.com/zhouhaoyi/ETDataset datasets (ETTm1, ETTm2, ETTh1, ETTh2). Additionally, we utilize 4 public datasets to assess the performance of short-term time series forecasting, which include: PEMS [^7] http://pems.dot.ca.gov (PEMS03, PEMS04, PEMS07, and PEMS08). These datasets have been widely used for benchmarking and are publicly available in [38], [41], [20], and [21]. Detailed dataset information is summarized in Table 1.
@@
Datasets Dim Horizon Frequency Dataset size Information Tasks
Weather 21 {96, 192, 336, 720} 10 min (36887,5270,10539) Weather Long-term forecasting
Electricity 321 {96, 192, 336, 720} Hourly (18412,2632,5260) Electricity Long-term forecasting
Solar-Energy 137 {96, 192, 336, 720} 10 min (36792,5256,10512) Electricity Long-term forecasting
ETTh1, ETTh2 7 {96, 192, 336, 720} Hourly (8640,2880,2880) Temperature Long-term forecasting
ETTm1, ETTm2 7 {96, 192, 336, 720} 15 min (34560,11520,11520) Temperature Long-term forecasting
Traffic 862 {96, 192, 336, 720} Hourly (12280,1756,3508) Transportation Long-term forecasting
ILI 170 {24, 36, 48, 60} 7 days (676,97,193) Illness Long-term forecasting
PEMS03 358 12 5 min (15724,5241,5243) Transportation Short-term forecasting
PEMS04 307 12 5 min (10195,3398,3399) Transportation Short-term forecasting
PEMS07 883 12 5 min (16934,5644,5646) Transportation Short-term forecasting
PEMS08 170 12 5 min (10713,3571,3572) Transportation Short-term forecasting
@@
Transformer-based: Informer [44], Autoformer [38], Crossformer [42], FEDformer [45], PatchTST [26], and iTransformer [23].
@@
xLSTM-based: xLSTMTime [1], P-sLSTM [19], and xLSTM-Mixer [20].
@@
For long-term forecasting, the prediction horizons for the ILI dataset are set to $T\in\{24,36,48,60\}$, with StoxLSTM’s look-back window fixed at $L=96$. For all other datasets, the prediction horizons are $T\in\{96,192,336,720\}$, and the look-back length of StoxLSTM is set to $L=336$. For short-term forecasting, all models are set to $T=12$ and $L=96$. Baseline results and look-back length $L$ settings for long-term forecasting are primarily sourced from PatchTST [26], xLSTM-Mixer [20], and TimeMixer++ [35], supplemented by our own experiments where results are unavailable. Evaluation metrics for long-term forecasting include Mean Squared Error (MSE) and Mean Absolute Error (MAE), while short-term forecasting is assessed using MAE, Mean Absolute Percentage Error (MAPE), and Root Mean Squared Error (RMSE). Experiments are conducted on an NVIDIA RTX 3090 GPU (24GB) and 3 NVIDIA A800 GPUs (80GB).